OpenStudy (anonymous):

Will medal and fan!! Which of the following demonstrates the Distributive Property? 3(4a + 2) = 12a + 6 3(4a + 2) = 12a + 2 3(4a + 2) = 4a + 6 3(4a + 2) = 7a + 5

OpenStudy (cwrw238):

3(4a+ 2) = 3*4a + 3*2 = 12 + 6

OpenStudy (cwrw238):

* that's 12a + 6